# Thermal Comfort in Sight: Thermal Affordance and its Visual Assessment for Sustainable Streetscape Design

### v1

@@ Abstract
In response to climate change and urban heat island effects, enhancing human thermal comfort in cities is crucial for sustainable urban development. Traditional methods for investigating the urban thermal environment and corresponding human thermal comfort level are often resource intensive, inefficient, and limited in scope. To address these challenges, we (1) introduce the concept of thermal affordance, which represents the inherent capacity of a streetscape to influence human thermal comfort based on its visual and physical features; and (2) a method to evaluate it (visual assessment of thermal affordance — VATA), which combines street view imagery (SVI), online and in-filed surveys, and statistical learning algorithms. VATA extracts five categories of image features from SVI data and establishes 19 visual-perceptual indicators for streetscape visual assessment. Using a multi-task neural network and elastic net regression, we model their chained relationship to predict and comprehend thermal affordance for Singapore. VATA predictions are validated with field-investigated OTC data, providing a cost-effective and scalable method to assess the thermal comfort potential of urban streetscape. This framework can inform streetscape design to support sustainable, livable, and resilient urban environments.
@@ 1.1 Urban heat environment and outdoor thermal comfort (OTC)
The growing effect of urban heat island (UHI) has brought the thermal environment of cities and their adaptive strategies into the spotlight (Alcoforado and Andrade, 2008; Li et al., 2018; Takebayashi and Takebayashi, 2015; Lai et al., 2019; Tseliou et al., 2022). UHI is detrimental to the built environment – it drives extreme climate changes (Chen et al., 2022; Steensen et al., 2022), increases energy consumption for cooling (Errebai et al., 2022), and deteriorates health due to poor air quality and decreased thermal comfort (Tan et al., 2010).
Outdoor thermal comfort (OTC), the subjective evaluation of one’s satisfaction with urban thermal conditions, is a key area of study (Chen and Ng, 2012; Shooshtarian et al., 2020). Poor OTC that makes people feel hotter can raise skin temperatures, increasing the risk of heat-related illness (da Silva Espinoza et al., 2023; Kovats and Hajat, 2008), and can even affect mental health effects (Piracha and Chaudhary, 2022; Hwong et al., 2022). The growing concern around OTC has driven the need for effective OTC evaluation methods (Aghamolaei et al., 2022). However, traditional methods, which focus on field surveys and environmental measurements, remain costly and limited in scalability and precision (Johansson et al., 2014; Evola et al., 2020; Xiong et al., 2022). The research by Li et al. (2024) offers a promising example of high-resolution thermal climate mapping by integrating urban morphology and meteorological data. Nevertheless, it lacks detailed consideration of the street environment and does not incorporate survey or measurement validation of human comfort at the street level. These constraints, amid climate change’s growing impact, call for improved methods to assess urban-scale OTC, for fostering sustainable street designs that enhance thermal comfort and urban livability.
@@ 1.2 Visual assessment of streetscape quality based on street view imagery (SVI)
Street view imagery (SVI) is a crucial tool for assessing the quality of streetscapes and their social impacts in urban environments (Gilge, 2016; Biljecki and Ito, 2021; Qiu et al., 2022; Liang et al., 2023). The visual composition of streetscapes provides a wealth of multidimensional information, enabling researchers to understand the socioeconomic implications of built environments and guide improvements in urban design (Lee and Park, 2023). SVI, paired with computer vision and GIS, has become a standard method for large-scale streetscape analysis (Cheng et al., 2017; Kang et al., 2020; Larkin et al., 2021; Ki et al., 2023). It supports both objective analysis of streetscape elements, including image features (IF) such as geometry, color, composition, and texture, and captures subjective visual-perceptual indicators (VPI) such as visual comfort, enclosure, complexity, through online surveys (Naik et al., 2014; Xu et al., 2022; Zhao et al., 2023). IF refers to the objective, measurable aspects of an image (SVI in urban studies) that can be extracted and analyzed through computational techniques (Xiao and Quan, 2009; Nagata et al., 2020), while VPI, on the other hand, refers to the subjective, human-centered interpretations of streetscapes, capturing how people perceive and experience the urban environment (Kang et al., 2020; Ogawa et al., 2024; Ito et al., 2024a). Many studies have invited participants to perform pairwise comparisons or grading-based visual assessments of SVI to evaluate streetscape quality and subjective visual perception they evoke (Kang et al., 2020; Larkin et al., 2021; Ito et al., 2024b).
@@
As the urban thermal environment and corresponding thermal comfort are closely linked to the objective composition of streetscapes and subjective visual perceptions (Nikolopoulou, 2011; Klemm et al., 2015; Yilmaz et al., 2021b; Hu et al., 2024; Yu et al., 2024), SVI has proven to be an important instrument in this domain. For instance, Hu et al. (2024) uses 3D morphological data from SVI for street-level thermal mapping, while Yu et al. (2024) combines SVI features with urban social-physical data for seasonal thermal assessments. When it comes to using streetscape elements and visual perception for thermal comfort evaluation, Klemm et al. (2015) identifies three key ways street greenery enhances thermal comfort across different street types, based on SVI and micrometeorological measurements: lowering radiant temperatures, improving streetscape aesthetics, and positively influencing conscious perceptions of microclimate. Similarly, after the evaluation of streetscape design through satellite images and SVI, local climate conditions, and ENVI-met modelling data, (Yilmaz et al., 2021b) suggests that thermal comfort can be improved through open spaces and thoughtful planting choices. Therefore, comprehensive visual assessments of streetscape quality not only enhance aesthetics and functionality but also improve thermal comfort and urban liveability.
@@ 1.3 Research gaps and our response
The interrelation between thermal comfort and streetscape design constitutes a significant area of study (Jamei et al., 2016). Based on existing literature (Klemm et al., 2015; Kang et al., 2020; Yilmaz et al., 2021b; Hu et al., 2024; Yu et al., 2024), we posit that both the objective composition of streetscape elements and subjective visual assessment could serve as proxies for evaluating a streetscape’s capacity to influence human thermal comfort. However, key research gaps persist.
@@
Firstly, there is no clear concept for assessing the built environment’s objective properties that influence thermal comfort in streetscape design. Although many studies have explored associations between streetscape elements, outdoor thermal environment, and OTC (Labdaoui et al., 2021; Sadeghi and Bahadori, 2021; Zheng et al., 2024; Yu et al., 2024; Hu et al., 2024), a theoretical framework that explains these relationships remains underdeveloped. This limits the creation of a convincing framework for sustainable streetscape design that improves outdoor thermal comfort.
@@
To address this gap, we introduce the concept of thermal affordance, inspired by Gibson’s theory of environmental affordance (Gibson, 1977). Our novel concept highlights the potential of urban design to promote thermal comfort through streetscape element configurations, forming the foundation for a data-driven framework to analyse their impact on OTC.
@@
Secondly, the use of human visual assessment in evaluating streetscapes’s thermal conditions and their capability to impact thermal comfort is rarely studied. While some research has touched on the interaction between visual information from SVI and outdoor thermal environments (Klemm et al., 2015; Urban et al., 2022; Zhang et al., 2022; Yan et al., 2023; Hu et al., 2024), they have not applied human visual assessment to examine how streetscape element composition influences thermal comfort. Some studies predict thermal environments using SVI (Yu et al., 2024; Hu et al., 2024), but they overlook OTC and lack validation through field experiments. Therefore, a methodological approach incorporating SVI and human visual assessment is needed for urban-scale OTC evaluation, validated by in-field OTC measurements. Moreover, while it is widely accepted that street greenery improves OTC (Klemm et al., 2015; Jamei et al., 2016; Motie et al., 2023), the influence of various streetscape elements on perceived thermal comfort remains underexplored, especially regarding buildings, roads, and vehicles.
@@
Our proposed visual assessment of thermal affordance (VATA) framework fills this gap by using SVI data and online questionnaires to quantify thermal affordance by integrating image features (IF) and human visual-perceptural indicators (VPI) from online surveys. VATA provides a novel approach to urban OTC evaluation.
@@
Third, the systematic application of multi-source urban data, such as SVI, human visual-perceptual surveys, and in-field comfort and physiological indicators, has not been fully adopted as an efficient workflow for assessing OTC in urban environments. Although some studies have linked SVI features with the urban thermal environment (Urban et al., 2022; Yan et al., 2022; Kim et al., 2023; Hu et al., 2024; Klimenko et al., 2024), a verifiable, replicable, and street-level methodology for urban-scale OTC evaluation supported by multi-source urban data has not yet been built.
@@ 1.4 Research question
2 Background and Related Work
2.1 The concept of OTC and traditional measurement approach
Thermal comfort is a psychological state where individuals express satisfaction with their surrounding thermal environment, encapsulating subjective perceptions of thermal sensations (ANSI/ASHRAE Standard 55, 2023). Six key factors influence thermal comfort: air temperature, radiant temperature, humidity, air speed, clothing insulation, and metabolic rate (ANSI/ASHRAE Standard 55, 2023).
Thermal comfort studies in the built environment are classified into indoor thermal comfort (ITC) and outdoor thermal comfort (OTC). ITC studies typically collect data from controlled climatic chambers or adaptive on-site surveys (Fanger et al., 1970; Cheng et al., 2012), examining changes in comfort in regulated or dynamic indoor environments (De Dear et al., 1998; Nicol and Humphreys, 2002; Lei et al., 2024b). Both methods contributed to the development of comprehensive thermal comfort indices, such as the new effective temperature (ET) and the standard effective temperature (SET), as proposed by Gagge (1971). Another model developed by Fanger et al. (1970) is the predicted mean vote (PMV), which predicts the average thermal sensation on a scale from -3 (cold) to +3 (hot), based on six environmental factors: air temperature, radiant temperature, relative humidity, air speed, clothing insulation, and metabolic rate. Recent ITC research has expanded these dimensions into active learning for thermal comfort modelling and improved personalised HVAC controls (Abdelrahman et al., 2022; Tekler et al., 2023; Ono et al., 2022), forming a comprehensive ITC improvement framework.
When it comes to OTC, urban environments, characterised by their intricate and complex nature, deviate significantly from the finite conditions of ITC evaluation. The multitude of diverse variables present in urban environments, such as solar radiation and shade availability, complicates climatic parameter control in OTC evaluations (Spagnolo and De Dear, 2003). This complexity hinders comprehensive OTC assessments in urban areas (Honjo, 2009). To address this challenge, Mayer and Höppe (1987) developed the physiological equivalent temperature (PET) for estimating outdoor thermal conditions comparable to indoor environments. SET, PMV, and PET are now widely used for OTC evaluation, though the complexity of urban meteorological variables and human factors makes the process challenging (Honjo, 2009).
Traditional OTC measurements rely on meteorological instruments, field investigations, and surveys (Acero and Herranz-Pascual, 2015; Mirzaei, 2015), but these methods are resource-intensive, limited in spatial coverage, and lack geographical precision (Aghamolaei et al., 2022). Such limitations highlight the need for a scalable, cost-effective framework for urban-scale OTC assessment that integrates multidimensional data while ensuring precision and applicability in sustainable urban planning.
2.2 Affordance as a potential link between human perception, behaviour and urban visual information encoded in SVI
2.2.1 The theory of affordance
The theory of affordance, proposed by Gibson (1977), highlights how the ecological environment shapes perception and behaviour. Applied to the built environment, affordances represent environmental features that promote human perception and behaviour (Gibson, 1977; Shashank et al., 2022). According to Gibson (1977), the environment acts as a ‘surface’ medium through which humans perceive and interact with their surroundings, offering clues to potential actions. The informational content and environmental characteristics embedded within the ’surfaces’ constitute the ‘value’ and ‘meaning’ of a place. These features, defined as affordances and inferred from environmental attributes, indicate likely human perceptions and behaviours even without direct experience. Gibson (1977) further emphasises that this inherent affordance of the environment can be deduced through visual assessment.
Affordance theory has been employed to elucidate the relationship between the built environment and human mobility behaviour, utilising geospatial indicators such as greenery index, street network morphology metrics, and urban function density, to comprehend how environments support activities like walking, biking, and running (Fusco, 2016; Shashank et al., 2022; Ito et al., 2024b). Public space density and green infrastructure density have been shown to be a significant factor in promoting walkability and liveability in cities (Littke, 2015; Zuniga-Teran et al., 2019). These relationships are supported by both qualitative studies (e.g., surveys and field observations) (Nissen et al., 2020) and quantitative analyses, where GIS data and human behaviour patterns show significant correlations in multivariate models (Kruse et al., 2021).
2.2.2 Affordance information from SVI in urban studies
SVI has become a key tool in urban studies, particularly for evaluating urban environment quality (Biljecki and Ito, 2021). As the built environment is designed to serve society, its affordances must be considered when conducting such evaluations. The built environment provides a foundation for urban life, influencing social activities, events, human movement, and exploration. Understanding these affordances is crucial for improving city conditions. SVI, as a direct and human-centred data source, offers valuable insights into the built environment’s affordances. Until now, in conjunction with SVI, affordance theory has been used to explain the evaluations of playability (Kruse et al., 2021), runnability (Shashank et al., 2022), and spatial navigation (Gregorians and Spiers, 2022), which has been shown to be a robust theoretical framework for explaining the relationship between urban attributes and human perception or behaviour (Shashank et al., 2022).
SVI provides rich visual data on urban streetscapes, capturing a variety of physical features that can be extracted using advanced computer vision techniques. Algorithms like PSPNet (Zhao et al., 2017) and FCN (Long et al., 2015), trained on datasets such as Cityscapes (Cordts et al., 2016) and ADE20K (Zhou et al., 2017b), enable precise semantic segmentation of SVI. These algorithms allow researchers to quantify the proportions of elements such as buildings, trees, and roads. These compositions offer insights into the affordances of urban streetscapes. For example, greenery indices derived from these images link higher green views to increased pedestrian activity and improved mental well-being (Wang et al., 2019a; He et al., 2020), while visual complexity affects perceived safety and comfort (Kawshalya et al., 2022).
In addition to streetscape indices based on element percentages, other metrics have been developed to capture additional dimensions of urban streetscapes in SVI. These include vehicle and pedestrian counts (Goel et al., 2018), the architectural style and building characteristics (Kang et al., 2018; Xu et al., 2023; Lei et al., 2024a), urban scene types (Su et al., 2021), and weather conditions (Ibrahim et al., 2019). These metrics, derived from specialised computer models and datasets, provide valuable spatial data for assessing urban environment affordances.
Beyond objective indices derived from computer vision and SVI datasets, human visual assessments of streetscapes involve subjective dimensions such as imageability, enclosure, and complexity, which are often interrelated (Qiu et al., 2022; Yang et al., 2023). These subjective qualities can be evaluated using SVI-based visual-perceptual questionnaires (Ogawa et al., 2024), typically conducted online. Participants rate or compare SVIs based on classic subjective indicators such as imageability, enclosure, human scale, transparency, complexity, and safety, reflecting their perceptions of urban spaces (Ewing and Handy, 2009; Qiu et al., 2022; Xu et al., 2022).
By combining objective SVI analysis with subjective assessment surveys, we can explore various urban spaces and their perceptual and behavioural potentials across different urban scales (Wang et al., 2023; Yang et al., 2023). This affordance analysis framework, using both objective and subjective metrics, effectively evaluates urban environments and streetscape design (Song et al., 2022; Qiu et al., 2023). These assessments can reveal a variety of human behaviours and urban outcomes, including walking and cycling patterns (Ito and Biljecki, 2021), property values (Qiu et al., 2023; Yang et al., 2023), and public health indicators (Larkin and Hystad, 2019; Wang et al., 2019b). Superior streetscape design improves the visual experience, walkability, bikeability, environmental appeal, and public activities, boosting the value of urban streets (Park et al., 2014; Gan et al., 2021; Strath et al., 2012).

### v4

@@ Abstract
{textblock*}
(0cm,-12cm)
This is the Accepted Manuscript version of an article published by Elsevier in the journal Building and Environment in 2025, which is available at: https://doi.org/10.1016/j.buildenv.2025.112569 Cite as: Yang S, Chong A, Liu P, Biljecki F (2025): Thermal comfort in sight: Thermal affordance and its visual assessment for sustainable streetscape design. Building and Environment, 271: 112569.
{textblock*}
1.5(-0.8cm,15cm) © 2025, Elsevier. Licensed under the Creative Commons Attribution-NonCommercial-NoDerivatives 4.0 International (http://creativecommons.org/licenses/by-nc-nd/4.0/) In response to climate change and urban heat island effects, enhancing human thermal comfort in cities is crucial for sustainable urban development. Traditional methods for investigating the urban thermal environment and corresponding human thermal comfort level are often resource intensive, inefficient, and limited in scope. To address these challenges, we (1) introduce a new concept named thermal affordance, which formalizes the integrated inherent capacity of a streetscape to influence human thermal comfort based on its visual and physical features; and (2) an efficient method to evaluate it (visual assessment of thermal affordance — VATA), which combines street view imagery (SVI), online and in-field surveys, and statistical learning algorithms. VATA extracts five categories of image features from SVI data and establishes 19 visual-perceptual indicators for streetscape visual assessment. Using a multi-task neural network and elastic net regression, we model their chained relationship to predict and comprehend thermal affordance for Singapore. VATA predictions are validated with field-investigated OTC data, providing a cost-effective, scalable, and transferable method to assess the thermal comfort potential of urban streetscape. Moreover, we demonstrate its utility by generating a geospatially explicit mapping of thermal affordance, outlining a model update workflow for long-term urban-scale analysis, and implementing a two-stage prediction and inference approach (IF-VPI-VATA) to guide future streetscape improvements. This framework can inform streetscape design to support sustainable, liveable, and resilient urban environments.
@@ 1.1 Urban heat environment and outdoor thermal comfort (OTC)
The growing effect of urban heat island (UHI) has brought the thermal environment of cities and their adaptive strategies into the spotlight [1, 2, 3, 4, 5]. UHI is detrimental to the built environment – it drives extreme climate changes [6, 7], increases energy consumption for cooling [8], and deteriorates health due to poor air quality and decreased thermal comfort [9].
Outdoor thermal comfort (OTC), the subjective evaluation of one’s satisfaction with urban thermal conditions, is a key area of study [10, 11]. Poor OTC that makes people feel hotter can raise skin temperatures, increasing the risk of heat-related illness [12, 13], and can even affect mental health [14, 15]. The growing concern around OTC has driven the need for effective OTC evaluation methods [16]. However, traditional methods, which focus on field surveys and environmental measurements, remain costly and limited in scalability and precision [17, 18, 19]. The research by [20] offers a promising example of high-resolution thermal climate mapping by integrating urban morphology and meteorological data. Nevertheless, it lacks detailed consideration of the street environment and does not incorporate survey or measurement validation of human comfort at the street level. These constraints, amid climate change’s growing impact, call for improved methods to assess urban-scale OTC, for fostering sustainable street designs that enhance thermal comfort and urban livability.
Urban environments, characterised by their intricate and complex nature, deviate significantly from the finite conditions of indoor thermal comfort (ITC) evaluation. The multitude of diverse variables present in urban environments, such as solar radiation and shade availability, complicates climatic parameter control in OTC evaluations [21]. This complexity hinders comprehensive OTC assessments in urban areas [22]. To address this challenge, [23] developed the physiological equivalent temperature (PET) for estimating outdoor thermal conditions comparable to indoor environments. Another model applied for OTC is the predicted mean vote (PMV), developed by [24], which predicts the average thermal sensation on a scale from -3 (cold) to +3 (hot), based on six environmental factors: air temperature, radiant temperature, relative humidity, air speed, clothing insulation, and metabolic rate. Recently, the universal thermal climate index (UTCI) has gained attention for its dynamic consideration of human physiological responses, making it particularly suitable for urban environments [25, 26]. PET, PMV, and UTCI are now widely used for OTC evaluation, though the complexity of urban meteorological variables and human factors makes the process challenging [22].
Traditional OTC measurements based on these indexes rely on meteorological instruments, field investigations, and surveys [27, 28], but these methods are resource-intensive, limited in spatial coverage, and lack geographical precision [16]. Such limitations highlight the need for a scalable, cost-effective framework for urban-scale OTC assessment that integrates multidimensional data while ensuring precision and applicability in sustainable urban planning. Moreover, recent advances in ITC research have introduced active learning approaches to thermal comfort modelling, enabling more adaptive and data-efficient solutions [29, 30, 31]. These developments, coupled with the growing demand for responsive planning frameworks, have paved the way for enhanced thermal comfort assessment systems that not only integrate changing environmental factors but also align with dynamic urban planning needs [32, 33]. Together, these trends underline the shift toward a more efficient, responsive and adaptive framework for urban-scale OTC evaluation, which is essential for sustainable streetscape design and human-centric urban development in the digital era.
@@ 1.2 Visual assessment of streetscape quality based on street view imagery (SVI)
Street view imagery (SVI) is a crucial tool for assessing the quality of streetscapes and their social impacts in urban environments [34, 35, 36, 37, 38, 39]. The visual composition of streetscapes provides a wealth of multidimensional information, enabling researchers to understand the socioeconomic implications of built environments and guide improvements in urban design [40]. SVI, paired with computer vision and GIS, has become a standard method for large-scale streetscape analysis [41, 42, 43, 44, 45, 46]. It supports both objective analysis of streetscape elements, including image features (IF) such as geometry, colour, composition, and texture, and captures subjective visual-perceptual indicators (VPI) such as visual comfort, enclosure, complexity, through online surveys [47, 48, 49]. IF refers to the objective, measurable aspects of an image (SVI in urban studies) that can be extracted and analysed through computational techniques [50, 51, 52, 53, 54], while VPI, on the other hand, refers to the subjective, human-centred interpretations of streetscapes, capturing how people perceive and experience the urban environment [42, 55, 56].
SVI captures a variety of IFs that can be extracted using advanced computer vision techniques. Algorithms like PSPNet [57] and FCN [58], trained on datasets such as Cityscapes [59] and ADE20K [60], enable precise semantic segmentation of SVI. These algorithms allow researchers to quantify the proportions of elements such as buildings, trees, and roads, and open-source packages are developed for semantic SVI data analytics in urban planning [61, 62, 63]. In addition to streetscape indices based on element percentages, other metrics have been developed to capture additional IFs of urban streetscapes in SVI. These include vehicle and pedestrian counts [64], the architectural style and building characteristics [65, 66, 67], urban scene types [68], and weather conditions [69]. These metrics, derived from specialised computer models and datasets, provide valuable spatial data for assessing urban environment.
Beyond objective IFs derived from computer vision and SVI datasets, human visual assessments of streetscapes involve subjective VPIs, which are often interrelated [36, 70]. These subjective qualities can be evaluated using SVI-based VPI questionnaires [55], typically conducted online. Participants rate or compare SVIs based on classic subjective VPIs such as imageability, enclosure, human scale, transparency, complexity, and safety, reflecting their perceptions of urban spaces [71, 36, 48]. Many studies have invited participants to perform pairwise comparisons or grading-based (such as score range 1-10) visual assessments of SVI to evaluate streetscape quality and subjective visual perception they evoke [42, 43, 72].
@@
As the urban thermal environment and corresponding thermal comfort are closely linked to the objective composition of streetscapes and subjective visual perceptions [73, 74, 75, 76, 77], SVI has proven to be an important instrument in this domain. For instance, [76] uses 3D morphological data from SVI for street-level thermal mapping, while [77] combines SVI features with urban social-physical data for seasonal thermal assessments. When it comes to using streetscape elements and visual perception for thermal comfort evaluation, [74] identifies three key ways street greenery enhances thermal comfort across different street types, based on SVI and micrometeorological measurements: lowering radiant temperatures, improving streetscape aesthetics, and positively influencing conscious perceptions of microclimate. Similarly, after the evaluation of streetscape design through satellite images and SVI, local climate conditions, and ENVI-met modelling data, [75] suggests that thermal comfort can be improved through open spaces and thoughtful planting choices. Therefore, comprehensive visual assessments of streetscape quality not only enhance aesthetics and functionality but also improve thermal comfort and urban liveability.
1.3 Affordance information from SVI in urban studies
The theory of affordance, proposed by [78], highlights how the ecological environment shapes perception and behaviour. Applied to the built environment, affordances represent environmental features that promote human perception and behaviour [78, 79]. According to [78], the environment acts as a ‘surface’ medium through which humans perceive and interact with their surroundings, offering clues to potential actions. The informational content and environmental characteristics embedded within the ’surfaces’ constitute the ‘value’ and ‘meaning’ of a place. These features, defined as affordances and inferred from environmental attributes, indicate likely human perceptions and behaviours even without direct experience. [78] further emphasises that this inherent affordance of the environment can be deduced through visual assessment.
Affordance theory has been employed to elucidate the relationship between the built environment and human mobility behaviour, utilising geospatial indicators such as greenery index, street network morphology metrics, and urban function density, to comprehend how environments support activities like walking, biking, and running [80, 79, 72, 81]. The built environment provides a foundation for urban life, influencing social activities, events, human movement, and exploration. Understanding these affordances is crucial for improving city conditions. SVI, as a direct and human-centred data source, offers valuable insights into the built environment’s affordances. Until now, in conjunction with SVI, affordance theory has been used to explain the evaluations of playability [82], runnability [79], and spatial navigation [83], which has been shown to be a robust theoretical framework for explaining the relationship between urban attributes and human perception or behaviour [79].
Both objective SVI indices and its subjective VPIs derived from human visual assessment offer insights into the affordances of urban streetscapes. For example, greenery indices derived from these images link higher green views to increased pedestrian activity and improved mental well-being [84, 85], while visual complexity affects perceived safety and comfort [86]. By combining objective SVI analysis with subjective SVI assessment surveys, we can explore various urban spaces and their perceptual and behavioural potentials across different urban scales [87, 70]. This affordance analysis framework, using both objective and subjective metrics, effectively evaluates urban environments and streetscape design [88, 89]. These assessments can reveal a variety of human behaviours and urban outcomes, including walking and cycling patterns [90], property values [89, 70], and public health indicators [91, 92]. Superior streetscape design improves the visual experience, walkability, bikeability, environmental appeal, and public activities, boosting the value of urban streets [93, 94, 95].
@@ 1.4 Research gaps and our response
The interrelation between thermal comfort and streetscape design constitutes a significant area of study [96]. Based on existing literature [74, 42, 75, 76, 77], we posit that both the objective composition of streetscape elements and subjective visual assessment could serve as proxies for evaluating a streetscape’s capacity to influence human thermal comfort. However, key research gaps persist.
@@
Firstly, there is no clear concept for assessing the built environment’s objective properties that influence thermal comfort in streetscape design. Although many studies have explored associations between streetscape elements, outdoor thermal environment, and OTC [97, 98, 99, 77, 76], a theoretical framework that explains these relationships remains underdeveloped. This limits the creation of a convincing framework for sustainable streetscape design that improves outdoor thermal comfort.
@@
To address this gap, we introduce the concept of thermal affordance (see details in Section 2), inspired by Gibson’s theory of environmental affordance [78]. Our novel concept highlights the potential of urban design to promote thermal comfort through streetscape element configurations, forming the foundation for a data-driven framework to analyse their impact on OTC.
@@
Secondly, the use of human visual assessment in evaluating streetscapes’ thermal conditions and their capability to impact thermal comfort is rarely studied. While some research has touched on the interaction between visual information from SVI and outdoor thermal environments [74, 100, 101, 102, 76], they have not applied human visual assessment to examine how streetscape element composition influences thermal comfort. Some studies predict thermal environments using SVI [77, 76], but they overlook OTC and lack validation through field experiments. Therefore, a methodological approach incorporating SVI and human visual assessment is needed for urban-scale OTC evaluation, validated by in-field OTC measurements. Moreover, while it is widely accepted that street greenery improves OTC [74, 96, 103], the influence of various streetscape elements on perceived thermal comfort remains underexplored, especially regarding buildings, roads, and vehicles.
@@
Our proposed visual assessment of thermal affordance (VATA) framework fills this gap by using SVI data and online questionnaires to quantify thermal affordance by integrating image features (IF) and human visual-perceptual indicators (VPI) from online surveys. VATA provides a novel approach to urban OTC evaluation.
@@
Third, the systematic application of multi-source urban data, such as SVI, human visual-perceptual surveys, and in-field comfort and physiological indicators, has not been fully adopted as an efficient workflow for assessing OTC in urban environments. Although some studies have linked SVI features with the urban thermal environment [100, 104, 105, 76, 106], a verifiable, replicable, and street-level methodology for urban-scale OTC evaluation supported by multi-source urban data has not yet been built.
@@ 1.5 Research question
  4. 4.
    What analytical insights and urban planning frameworks can be derived from the data outputs of the VATA framework to guide sustainable streetscape design?